OpenStudy (anonymous):

An attractive force of 7.2 N occurs between two point charges that are 0.10 m apart. If one charge is -4.0 µC, what is the other charge? (µC = 1.0 × 10^-6 C) -4.0 µC , -2.0 µC , +2.0 µC , or +4.0 µC

OpenStudy (anonymous):

i think it's C.

OpenStudy (michele_laino):

here we have to apply teh law of Coulomb

OpenStudy (michele_laino):

the*

OpenStudy (anonymous):

Not so sure! sorry!

OpenStudy (michele_laino):

\[\Large F = K\frac{{{Q_1}{Q_2}}}{{{d^2}}}\]

OpenStudy (anonymous):

+2.0 µC

OpenStudy (anonymous):

okay! what do we plug in?

OpenStudy (michele_laino):

for example I will solve that fromula for Q2, and I get: \[{Q_2} = \frac{{F{d^2}}}{{K{Q_1}}}\]

OpenStudy (anonymous):

7.2 should be plugged in

OpenStudy (anonymous):

ohh what would that look like? :/ i am confused :/ is is 7.2 divided by somehing?

OpenStudy (michele_laino):

substituting our data, we can write: \[\Large {Q_2} = \frac{{F{d^2}}}{{K{Q_1}}} = \frac{{7.2 \times {{\left( {{{10}^{ - 1}}} \right)}^2}}}{{9 \times {{10}^9} \times \left( { 4 \times {{10}^{ - 6}}} \right)}}\]

OpenStudy (anonymous):

F = k*q1*q2/r^2 so q2 = F*r^2/(k*q1) = 7.2*0.10^2/(9.0x10^9*4.0x10^-6) = 2.0x10^-6 and since the force is attractive the charge must be C - 2.0 µC

OpenStudy (anonymous):

sorry b! not c!

OpenStudy (michele_laino):

no, it is C, since the force is attractive, and only opposite charges interact with an attractive force

OpenStudy (anonymous):

ohh so our solution is choice C?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay! thank you both!!

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

was that your last question?

OpenStudy (anonymous):

oh never mind